a spacecraft of mass 100 kg breaks into two when its velocity is 10 m/sec. after the break, a mass of 10 kg of the space craft is left stationary. what is the velocity of remaining part?

Grade:8

1 Answers

Ameena
4 years ago
Before explosion,
mass of spacecraft = 100
velocity=10
momentum = 1000
 
After explosion ,
mass of one part = 10
velocity=0
mass of the other part = 100-10
velocity = v ?
momentum = 10-0+(100-10)v
                   = (100-10)v
 
conserving momentum,
100x10=(100-10)v
v=1000/100-10
therefore velocity of other part is 1000/90 = 11.11m/s